OpenStudy (anonymous):

I will award you! :) How many solutions does the system of equations have? -2=2x-4 y=-2/3x-1 A. 0 B. 1 C. 2 D. infinite

OpenStudy (anonymous):

Answer is F

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Are you sure you have written the system correctly ?

OpenStudy (anonymous):

i wrote it wrong. one sec. :/ haha

OpenStudy (anonymous):

2x+3y=6 y=-2/3x-1

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2x+3y=6 }\) \(\normalsize\color{blue}{ y=(-2/3)x-1 }\) just want to play with the first equation a little bit to show you something. \(\normalsize\color{blue}{ 2x+3y=6 }\) \(\normalsize\color{blue}{ 3y=-2x+6 }\) \(\normalsize\color{blue}{ y=(-2/3)x+2 }\) so what you have is the following system above in red. \(\normalsize\color{red}{ y=(-2/3)x+2 }\) \(\normalsize\color{red}{ y=(-2/3)x-1 }\)

OpenStudy (solomonzelman):

how many solutions can here be ?

OpenStudy (anonymous):

two?

OpenStudy (solomonzelman):

Okay, lets do this a little differently. When I say \(\normalsize\color{red}{ y=(-2/3)x-1 }\) \(\normalsize\color{red}{ y=(-2/3)x+2 }\) and when I say \(\normalsize\color{blue}{ y+(2/3)x=-1 }\) \(\normalsize\color{blue}{ y+(2/3)x=2 }\) I am not changing anything, right?

OpenStudy (anonymous):

right.

OpenStudy (solomonzelman):

Now, here the problem arises. In the first equation ( \(\normalsize\color{blue}{ y+(2/3)x=-1 }\) ) we see that `y+(2/3)x` is equal to `1` but in the second equation we see that this same number (which is `y+(2/3)x` ) is equal to 2. Can the same number be equal to -1 and to 2 at the same time ?

OpenStudy (anonymous):

absolutely not.

OpenStudy (solomonzelman):

correct, and that means that there are no solutions.... correct:)

OpenStudy (solomonzelman):

no solutions is same as zero solutions.

OpenStudy (anonymous):

thank you so much! you helped me out so much just in the nic of time! :)

OpenStudy (solomonzelman):

Anytime! Btw, where do you live? I can guess somewhere south from the equator since you have school, but where if not a secret? (just wondering)

OpenStudy (anonymous):

Pennsylvania. summer school.. :/

OpenStudy (solomonzelman):

Ohh... I see:) I live in Atlanta GA (absolute silence and good weather) anyway, good luck with your summer school !

OpenStudy (anonymous):

Ah! i love Georgia. well, thanks again and have a good summer! :)

OpenStudy (anonymous):

I'm fairly certain that i'll call upon you again for help.

OpenStudy (solomonzelman):

You too: good summer:) YOu are probably finishing your summer school right now, so have a good real summer, after your school summer.

OpenStudy (anonymous):

haha, not quite yet.. i finish in about one month.. by which point i'm starting regular school. :(